OpenStudy (anonymous):

Only 1 question. I need help on it Has 2 parts! Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 5x + 3y = 8 y = −2x + 1

OpenStudy (anonymous):

Substitution is really nice in cases where it's very easy to solve one of the equations for one of the variables. Elimination tends to be a great choice whenever you can look at the equations and notice that one of the terms will cancel out really nicely. Graphing is just... kind of lame. It's good as a mental concept to have, but as an actual method of solution, it tends to be more work than it is worth. However, if they are very easy to graph, then I guess it could be a good choice.

OpenStudy (anonymous):

Thank you :) I was going too say substitution it is the easiest :D

OpenStudy (anonymous):

I agree =) Your first step for substitution method would be to pick an equation and pick a variable, then solve that equation for that variable. But for this one, hey look! That's already done for me. I have y=-2x+1

OpenStudy (precal):

if you like to graph and have a graphing calculator, then that is also, an easier method.

OpenStudy (anonymous):

If you have a graphing calculator, sure. That fixes most of the drawbacks of the method. You don't have to draw a scale. You don't get the imperfect lines that you would get by hand drawing. You don't have to look and guess what the point is that they are crossing because you can just use your calculator's intersection tool. The only remaining difficulty I would say is the viewing window.

OpenStudy (anonymous):

I tried the Graphing cal and was a little more difficult for me.

OpenStudy (anonymous):

Then go with substitution =) And I'd be happy to check your answer for the second part if you type it up here.

OpenStudy (anonymous):

Okay :) hold on let me do it real quick!

OpenStudy (precal):

is it a ti-84 or ti-83? if so 2nd trace 5, enter, enter, enter

OpenStudy (precal):

(-5,11)

OpenStudy (anonymous):

Giving the answer there, totally not necessary.

OpenStudy (anonymous):

5x + 3y = 8 2x + y = 1 -x + 0y = 5 -x = 5 x = -5 5 (-5) + 3y = 8 -25 + 3y = 8 3y = 8 + 25 3y = 33 y = 11. this is how i did it earlier....

OpenStudy (anonymous):

You realize that is the elimination method instead of the substitution method?

OpenStudy (anonymous):

lol really? my teacher started me off like that ... i didnt know which one it was.

OpenStudy (anonymous):

Your work all looks fine though.

OpenStudy (anonymous):

You said that substitution would be easiest. Why did you say that?

OpenStudy (anonymous):

Actually, I have to go eat dinner. Sorry!

OpenStudy (anonymous):

because i thought that was substitution..

OpenStudy (precal):

that's ok I can help at this point

OpenStudy (anonymous):

No, that is elimination. Substitution would be when I solve one equation for one variable and then I SUBSTITUTE for that variable in the OTHER equation.

OpenStudy (anonymous):

Go get 'em precal.

OpenStudy (precal):

You go have a nice dinner, I have this

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

Thank you both :} my teacher didn't really explain it that well for me to understand it.

OpenStudy (precal):

let me check something real quick

OpenStudy (anonymous):

okay :)

OpenStudy (precal):

I am going to delete my slides, don't go anywhere I will redo them

OpenStudy (anonymous):

okay that is fine :) i wanna see how to work them .

OpenStudy (precal):